OpenStudy (anonymous):

-12a^-5/2a^-8 simplify do not use negative exponents in the answer please help get me started

OpenStudy (anonymous):

When you have negative exponents, you put the thing with the negative exponent into the denominator, with what ever is left on top. Example: \[45x ^{-2}= \frac{ 45 }{ x ^{2} }\] Now if you have fractions, its more of a multi step problem. Its going to be a complex fraction, with one fraction divided by another. A big mistake people make is that they just flip the negatives from the top to the bottom or vice versa. Heres an example: \[\frac{12p ^{3}q ^{-1} }{ 8p ^{-2}q }=\frac{ \frac{ 12p ^{2} }{ q } }{ \frac{ 8q }{ p ^{2} } }\]

OpenStudy (anonymous):

Let me know when you get it and want the next step

OpenStudy (anonymous):

I am so lost at fractions to begin with then they throw in all these damn letters I'm thinking from what I am reading in my book is that I multiply 12 5's

OpenStudy (anonymous):

It's ok, ill help ya :P Nope, you dont multiply that. Do you get what you do if you have a negative exponent, justa simple one?

OpenStudy (anonymous):

i know that when you add two negatives they equal a positive, etc lil stuff like that

OpenStudy (anonymous):

See the first part that i showed you? Do you get that?

OpenStudy (anonymous):

i'm sorry i had to go pick up my son from school

OpenStudy (anonymous):

Oh ok.. thats fine

OpenStudy (anonymous):

i'm going to have to get back on a lil later thank you for your help though